# Optimal Task Assignment and Path Planning using Conflict-Based Search with Precedence and Temporal Constraints

### v2

@@ 2.3. Multi-Agent Reinforcement Learning (MARL)
MARL is a branch of RL whose objective is to train agents who can learn to cooperate or compete with each other in complex environments to achieve a common or individual goal. In partially observable Markov game Littman (1994) of $N$ agents, the set of states, actions and observations are given by $\mathcal{S}$, $\mathcal{A}_{1},\dots,\mathcal{A}_{N}$ and $\mathcal{O}_{1},\dots,\mathcal{O}_{N}$ respectively for each agent. A distribution $\rho:\mathcal{S}\xrightarrow[]{}[0,1]$ selects the initial states. Each agent obtains an observation, $o_{i}:\mathcal{S}\xrightarrow[]{}\mathcal{O}_{i}$, follows a stochastic policy, $\bm{\pi}_{\theta_{i}}:\mathcal{O}_{i}\times\mathcal{A}_{i}\xrightarrow[]{}[0,1]$ that generates the next state following the state transition function, $\mathcal{T}:\mathcal{S}\times\mathcal{A}_{1}\times\dots\times\mathcal{A}_{N}\xrightarrow[]{}\mathcal{S}$ and receives a reward, $r_{i}:\mathcal{S}\times\mathcal{A}_{i}\xrightarrow[]{}\mathbb{R}$. Each agent aims to maximize its own total expected return, $G_{t}=\sum^{T_{term}}_{k}\gamma^{k}r_{t+k+1}$, which is the total discounted reward from timestep $t$, where $\gamma$ and $T_{term}$ are the discount rate and the terminal timestep respectively.
@@ 5.2.2. Algorithm Setup
For comparisons with MAPF baselines, we implement a variant of CBS-TA Hönig et al. (2018) that maximizes return under the same subtask assignment framework as CBS-TA-PTC highlighted in Algorithm 1 but with vertex/edge conflicts only. Under CBS-TA, upon the explosion of a bomb at a given timestep, CT nodes are generated in a naive manner with vertex conflicts for each agent for that timestep, i.e. at least one agent is doing something wrong at that timestep. This variant of CBS-TA can be seen as CBS-TA-PTC without absolute temporal range, inter-goal temporal and precedence conflicts, It attempts to solve the TAPF-PTC problem by relying solely on vertex conflicts to resolve the underlying precedence and temporal conflicts through a more exhaustive and inefficient best-first search through the CT with more CT nodes relative to CBS-TA-PTC. Nevertheless, it is worth highlighting that this variant of CBS-TA is optimal and complete like CBS-TA-PTC as elaborated in Section 4.3, under the assumption that the entire task is a single subtask.
@@ 5.3.2. CBS
From Figure 2(c), the success rate and optimality ratio naturally decrease with an increasing seconds per timestep parameter for both CBS-TA-PTC and CBS-TA given the growing difficulty of the task with shorter bomb fuses and countdowns in terms of timesteps. By calculating the average inter-node distance and using the default walking speed of $4.317\>ms^{-1}$ in Minecraft, we find that $2$ seconds per timestep is the best approximation of the Minecraft dynamics, to which we have a success rate of approximately $0.65$ for CBS-TA-PTC compared to $0.15$ for CBS-TA. Hence, it is clear that the usage of vertex and edge conflicts alone in CBS-TA is insufficient to solve TAPF-PTC problems efficiently in general.

### v3

@@ 2.3. Multi-Agent Reinforcement Learning (MARL)
MARL is a branch of RL whose objective is to train agents who can learn to cooperate or compete with each other in complex environments to achieve a common or individual goal. In partially observable Markov game Littman (1994) of $N$ agents, the set of states, actions and observations are given by $\mathcal{S}$, $\mathcal{A}_{1},\dots,\mathcal{A}_{N}$ and $\mathcal{O}_{1},\dots,\mathcal{O}_{N}$ respectively for each agent. A distribution $\rho:\mathcal{S}\xrightarrow[]{}[0,1]$ selects the initial states. Each agent obtains an observation, $o_{i}:\mathcal{S}\xrightarrow[]{}\mathcal{O}_{i}$, follows a policy, $\bm{\pi}_{\theta_{i}}:\mathcal{O}_{i}\times\mathcal{A}_{i}\xrightarrow[]{}[0,1]$ that generates the next state following the state transition function, $\mathcal{T}:\mathcal{S}\times\mathcal{A}_{1}\times\dots\times\mathcal{A}_{N}\xrightarrow[]{}\mathcal{S}$ and receives a reward, $r_{i}:\mathcal{S}\times\mathcal{A}_{1}\times\dots\times\mathcal{A}_{N}\xrightarrow[]{}\mathbb{R}$. Each agent aims to maximize its expected return $\mathbb{E}(G_{i}^{t})$, where return $G_{i}^{t}=\sum^{T_{term}}_{k=0}\gamma^{k}r_{i,t+k+1}$, is the sum of discounted rewards from timestep $t$ and $\gamma$ and $T_{term}$ are the discount rate and the terminal timestep respectively.
@@ 5.2.2. Algorithm Setup
For comparisons with MAPF baselines, we implement a variant of CBS-TA Hönig et al. (2018) that maximizes return under the same subtask assignment framework as CBS-TA-PTC highlighted in Algorithm 1 but with vertex/edge constraints only. Under CBS-TA, upon the explosion of a bomb at a given timestep, CT nodes are generated in a naive manner with vertex constraints for each agent for that timestep, i.e. at least one agent is doing something wrong at that timestep. This variant of CBS-TA can be seen as CBS-TA-PTC without absolute temporal range, inter-goal temporal and precedence conflicts, It attempts to solve the TAPF-PTC problem by relying solely on vertex constraints to resolve the underlying precedence and temporal conflicts through a more exhaustive and inefficient best-first search through the CT with more CT nodes relative to CBS-TA-PTC. Nevertheless, it is worth highlighting that this variant of CBS-TA is optimal and complete like CBS-TA-PTC as elaborated in Section 4.3, under the assumption that the entire task is a single subtask.
@@ 5.3.2. CBS
From Figure 4(a), the success rate and optimality ratio naturally decrease with an increasing seconds per timestep parameter for both CBS-TA-PTC and CBS-TA given the growing difficulty of the task with shorter bomb fuses and countdowns in terms of timesteps. By calculating the average inter-node distance and using the default walking speed of $4.317\>ms^{-1}$ in Minecraft, we find that $2$ seconds per timestep is the best approximation of the Minecraft dynamics, to which we have a success rate of approximately $0.65$ for CBS-TA-PTC compared to $0.15$ for CBS-TA. Hence, it is clear that the usage of vertex and edge conflicts alone in CBS-TA is insufficient to solve TAPF-PTC problems efficiently in general.